# Funding AI for Good: A Call for Meaningful Engagement

### v1

@@ 5.2. Reframing AI4SG Funding: Recommendations for Funding Call Designs
Building on the more balanced approaches and existing gaps we identified in our dataset, we offer the following recommendations for funding call designs, putting them in conversation with existing literature on “tech for good” projects. We structure these recommendations around the four key components of funding calls, to serve as a template to inform the design of future funding calls.
@@ 5.2.1. Background: situating AI4SG in well-scoped contexts
Funding calls should orient towards projects that address well-scoped and contextually relevant problems. This doesn’t mean that funders should pre-define the problems, but they can support applicants by offering examples of relevant AI use cases or existing literature review on the social domain (Section 4.1.2). It may also involve requiring the active participation of communities during the problem formulation and solution brainstorming stage, where applicants and communities collaboratively investigate which technologies are appropriate for their specific problem. In a systematic review of community-collaborative approaches (CCA), Cooper et al. highlight that many projects involve communities only during limited stages (e.g., delivery of activities), rather than throughout the full research process—including problem framing and data analysis (Cooper et al., 2022). Yet, as Bratteteig and Wagner argue, engaging communities in deciding which problems to address is key, not only for inclusion but for redistributing power and ensuring they have real influence over project direction and outcomes (Bratteteig and Wagner, 2016). This level of participation reflects a deeper form of agency than simply evaluating or implementing pre-defined solutions (Delgado et al., 2023).
@@ 5.2.2. Expected Outcomes: defining success with intended communities
When real-world deployment is valued by the project teams and intended communities, it would be important to provide funding support and maintenance resources post-deployment—something notably absent from most funding programs in our dataset. In categorizing ICTD projects’ outcomes, Heeks calls attention to the high rate of projects that might have initially succeeded but break down over time (Heeks, 2002). As Jackson argues, “repair and maintenance” is a critical part of technology design, yet it pervasively falls out of view when the dominant discourse of technology focuses on innovations as a production, and not a process (Jackson, 2014). The maintenance work of a technology system is not only necessary for sustaining the system that will inevitably keep breaking (Jackson, 2014), it also builds relationships among workers who collaborate on the practical work (Irani and Silberman, 2014; D’Ignazio, 2024). Sustainable maintenance can involve building local capacity and intentionally transferring technology ownership to communities. For example, the research team behind SignSupport, a mobile app co-developed with the Deaf Community of Cape Town (DCCT) to support pharmacy interactions using South African Sign Language, trained community members to manage a computer lab and take on system administration and monitoring roles, while researchers stepped back to allow community leadership to grow without interference (Dearden and Tucker, 2016).
@@
Funding calls should be structured to include intended community members as co-leaders or co-applicants, ensuring they have the power to shape project direction from the outset—and importantly, that they receive funding directly. As Bratteteig and Wagner point out, the “skill composition” of a design team significantly shapes how problems and solutions are defined (Bratteteig and Wagner, 2016). Including community members in core project roles enables them to define the problem space and success criteria, and ensures their labor and expertise are properly acknowledged and compensated (Pine et al., 2020). This approach also helps redistribute power within AI4SG collaborations, supporting more equitable relationships between communities and research institutions (Erete et al., 2023). Crucially, when engaging communities throughout the project, it is important to recognize and compensate for the various forms of labor they contribute—from sharing lived experience and domain expertise to facilitating connections within their communities. Future research could work towards quantifying both visible and invisible labor (see (Toxtli et al., 2021) and (Li et al., 2022)), while also acknowledging that some forms of labor in community engagement, such as emotional labor, may be difficult or impossible to measure.
@@ 5.2.4. Funders’ Support: valuing community engagement as a necessary part of AI4SG work
Funders can better support community engagement in AI4SG by recognizing that relationship-building is a foundational phase of the work, which often occurs before a clear project direction is even established (Carroll and Rosson, 2013; Cooper et al., 2022). This phase is essential to community-collaborative approaches (Cooper et al., 2022), and it requires project teams’ routine presence that takes time and emotional labor (Dearden and Tucker, 2016). Long-term relationships allow shared agendas to emerge, with community priorities guiding technology development from the start (Dearden and Tucker, 2016; Carroll and Rosson, 2013). One concrete step funders can take is to extend the typical AI4SG funding timeline (in our corpus: with a median of one and a half years) to accommodate this early-stage work. Relationship-building alone can take months (Erete et al., 2023) or even years (Carroll and Rosson, 2013; Dearden and Tucker, 2016), and should be supported as a formal project phase, even prior to defining problems or solutions. One key challenge is determining appropriate timelines and support—whether financial or capacity-building—for this work. Future research in CCA could analyze researchers’ reflections (e.g., (Erete et al., 2023; Dearden and Tucker, 2016; Carroll and Rosson, 2013)) to better understand the duration and nature of relationship-building across projects.
@@
Furthermore, funders should also treat community engagement as a critical skill requiring training and support—on par with technical expertise. Many AI practitioners are trained in Computer Science departments that don’t routinely provide training in community engagement. As Okolo and Lin observe, AI4SG practitioners often use a hodgepodge of ad-hoc approaches to engage with end users (Okolo and Lin, 2024). Funders should recognize this knowledge gap and provide clear guidance for project teams to properly identify stakeholders or engage with end beneficiaries. There are multiple places where this guidance could be provided—as part of the funding call itself or additional training support that many funders already provide. Funders could consider citing existing guides in community engagement from the HCI community, such as the Equity-Centered Community Design Field Guide (Creative Reaction Lab, 2024), the Atlanta Community Engagement Playbook (Taranji Alvarado et al., 2024), and “readiness for partnership” checks (Pine et al., 2020). Other design toolkits like the Envisioning Cards (Friedman et al., 2024) might also be useful for AI4SG project teams to identify and engage with stakeholders. Since AI introduces additional hurdles to community-based designs, including requiring intensive data work (Miceli and Posada, 2022; Passi and Sengers, 2020; Sambasivan et al., 2021) and transparency issues (Okolo and Lin, 2024), HCI researchers could support funders by tailoring existing toolkits to the specific contexts of AI4SG, building on emerging Participatory AI4SG frameworks like the PACT framework (Bondi et al., 2021).
@@ 5.3. Calling for Meaningful Engagement: Recommendations for the Funding Design Process
Beyond the design of individual funding calls, we offer recommendations for funders on the design process of funding programs. The design process of grantmaking is influenced by funders’ connections to other stakeholders (Saha et al., 2022a). From our dataset, we observed that private tech companies were not only the providers of AI products, resources, and training support, but also AI4SG funders themselves. Through investment in AI4SG, private tech companies could not only expand their market influence (Henriksen and Richey, 2022; Lukacz, 2024; Nost and Colven, 2022) but also influence other funders’ beliefs about AI’s potential and their funding priorities. At the same time, the influence of communities on funding documents remains scant (c.f.  (Saha et al., 2022a)). To reorient AI4SG towards a more balanced approach, we see the need for funders to engage with intended communities and the HCI community.
5.3.1. Engaging with intended communities
Engaging with communities in the design of funding calls would enable funders to align their agendas with the needs and visions of those communities early in the decision-making process, before projects even begin (Saha et al., 2022a; Balestrini et al., 2014). For funding programs that target projects in a region that is far away from the funding organizations’ locations, it would be important to have regional partners with a more solid contextual understanding, such as regional research hubs and organizations. They can serve as “intermediaries” between funding organizations and intended end beneficiaries that funding organizations don’t have direct access to (Dell and Kumar, 2016; Saha et al., 2022a).
@@ 5.3.1. Engaging with intended communities
5.3.2. Engaging with the HCI community
Following up on Saha et al.’s call on HCI researchers and practitioners to serve as “mediators” between decision-makers and communities’ voices (Saha et al., 2022b), we see opportunities in the AI4SG space where HCI researchers and practitioners can mediate between communities, funders, and other stakeholders like technologists and funders’ regional partners. Specifically, the HCI community can support funders in identifying key stakeholders to engage with. HCI practitioners could also help funders design training materials on community engagement for project teams and provide community engagement workshops throughout the project stages. While there is a rich body of resources for community-collaborative approaches (CCA) (Cooper et al., 2022), there exist gaps for the HCI community to bridge between frameworks and practices, such as implementing formal evaluations of participatory design projects (Bossen et al., 2016; Nguyen, 2022).
@@ 5.3.2. Engaging with the HCI community
Moreover, the HCI community can provide tools and facilitations to foster reflexivity in the grantmaking process, such as reflection workshops that bring funders, regional partners, representatives of intended beneficiaries, and targeted applicants together to learn from past funding program experiences. Wellcome Trust’s design process of Mental Health Data Prizes (C8 and C9) was a good example of benefits of feedback workshops (Section 4.4.2). HCI researchers and practitioners can provide consultations on reflection workshops or deliberation tools when considering strategies around AI (e.g., (Kawakami et al., 2024b)).
@@
Last but not least, the HCI community can offer counter-narratives to the more techno-centric conceptualizations of AI in funding calls, assisting funders to consider alternatives that emphasize diverse ways of knowing and sustainable development of AI within planetary boundaries. Some emerging critical reassessments of current AI design practices and public discourse around AI include work on AI abandonment (Johnson et al., 2024), post-capitalist AI systems based on feminist values (Varon et al., 2024), Indigenous thoughts (Ofosu-Asare, 2024; Lewis et al., 2024), post-colonial computing (Sultana and Ahmed, 2019), and post-growth philosophies (Sharma et al., 2023; Kuijer and Laschke, 2024).
5.4. Limitations and Future Work
The list of funding documents in our dataset is neither exhaustive nor statistically representative. Our analysis is restricted to English-language materials and funders located in English-speaking regions, which excludes funding programs described in other languages. Additionally, funding programs that do not use the keywords in our study might have been overlooked. Future work could expand the analysis to AI4SG funding documents published in languages other than English. Furthermore, while calls for proposals and grant announcements reveal funders’ intentions, they do not necessarily reflect what projects actually get funded. Building on this work, future research could examine the characteristics of funded projects. This line of work could reveal alignment or gaps between vision and practice, including who receives funding (whether AI-capable organizations or mission-driven groups). Future work could also explicitly examine project outcomes to categorize commonalities of success and failures specific to AI4SG (similar to Dodson et al. (2012)’s analysis of ICTD projects). Since publicly available written materials may be biased toward success stories, focusing on one or a few case studies—with interviews with project teams and analysis of project-related documents—could yield insightful analysis. Finally, we emphasize that the funding documents in our corpus fall along a spectrum from more techno-centric to more balanced approaches, and we caution readers against viewing this spectrum as a dichotomy. Moreover, future work could further explore what it may look like to over-emphasize contextual understanding and under-emphasize technical capacity.
@@ 6. Conclusion
Artificial Intelligence for Social Good (AI4SG) emerges as a growing body of research and practice that explores the potential of AI and data-driven technologies to tackle social problems like public health and humanitarian aid (Shi et al., 2020; Perrault et al., 2020; Tomašev et al., 2020; Rolnick et al., 2022). However, a growing body of evidence suggests that AI4SG initiatives do not necessarily yield the intended positive outcomes in real-world contexts (Lin et al., 2024; Aula and Bowles, 2023; Ben Green, 2019). While prior work has focused on the actions and roles that project teams—typically AI developers and their nonprofit partners–play in project outcomes, an emerging body of work points to the significant influence of funders in shaping the direction, design, and priorities of technology initiatives aimed at social impact (Lin et al., 2024; Saha et al., 2022a; Gardner et al., 2022). While prior work has highlighted funders’ significant influence, we still lack a direct understanding of how funders frame AI’s role in addressing social good challenges and how they envision AI’s impact could be achieved.
In this study, we fill this gap and “study up” AI4SG funding documents by examining them through the lens of existing literature on “tech for good” projects. Our analysis reveals dissonances between AI4SG funders’ intentions for positive, sustainable social impact and the techno-centric approaches that some funding programs promoted. At the same time, we identified several more balanced approaches as well as existing gaps in funding calls. Building on our analysis, we offer recommendations for funders in funding call designs, while emphasizing the need for community engagement in the funding call design process. Finally, we call on the HCI community and community-collaborative approaches practitioners and researchers to mediate between AI4SG stakeholders to incorporate meaningful engagement with communities in the early stages of AI4SG projects—when funding programs are being designed.
@@ Appendix A Funding documents
C6 Data for Social Sustainability call for proposals 2021 Up to \$80,000 (avg) + AWS credits (up to \$20,000 avg) 1 year Amazon Private Tech Company United States Global human rights protections “full-time faculty member or permanent NGO researcher” (Amazon, 2021)

### v2

@@ 5.2. Reframing AI4SG Funding: Recommendations for Funding Call Designs
Building on the more balanced approaches and existing gaps we identified in our dataset, we offer the following recommendations for funding call designs, putting them in conversation with existing literature on “tech for good” projects. We structure these recommendations around the four key components of funding calls, to serve as a template to inform the design of future funding calls for funders, as well as future HCI research directions in Community-Collaborative Approaches (CCA).
@@ 5.2.1. Background: situating AI4SG in well-scoped contexts
Funding calls should orient towards projects that address well-scoped and contextually relevant problems. This doesn’t mean that funders should pre-define the problems, but they can support applicants by offering examples of relevant AI use cases or existing literature review on the social domain (Section 4.1.2). It may also involve requiring the active participation of communities during the problem formulation and solution brainstorming stage, where applicants and communities collaboratively investigate which technologies are appropriate for their specific problem. In a systematic review of CCA, Cooper et al. highlight that many projects involve communities only during limited stages (e.g., delivery of activities), rather than throughout the full research process—including problem framing and data analysis (Cooper et al., 2022). Yet, as Bratteteig and Wagner argue, engaging communities in deciding which problems to address is key, not only for inclusion but for redistributing power and ensuring they have real influence over project direction and outcomes (Bratteteig and Wagner, 2016). This level of participation reflects a deeper form of agency than simply evaluating or implementing pre-defined solutions (Delgado et al., 2023).
@@ 5.2.2. Expected Outcomes: defining success with intended communities
When real-world deployment is the goal, it is essential to provide funding support and maintenance resources post-deployment, which is notably absent from most funding programs in our dataset. In categorizing ICTD projects’ outcomes, Heeks calls attention to the high rate of projects that might have initially succeeded but break down over time (Heeks, 2002). As Jackson argues, “repair and maintenance” is a critical part of technology design, yet it pervasively falls out of view when the dominant discourse of technology focuses on innovations as a production, and not a process (Jackson, 2014). The maintenance work of a technology system is not only necessary for sustaining the system that will inevitably keep breaking (Jackson, 2014), it also builds relationships among workers who collaborate on the practical work (Irani and Silberman, 2014; D’Ignazio, 2024). Sustainable maintenance can involve building local capacity and intentionally transferring technology ownership to communities. For example, the research team behind SignSupport, a mobile app co-developed with the Deaf Community of Cape Town to support pharmacy interactions, trained community members to manage a computer lab and take on system administration roles, while researchers stepped back to allow community leadership to grow without interference (Dearden and Tucker, 2016).
@@
Funding calls should be structured to include intended community members as co-leaders or co-applicants, ensuring they have the power to shape project direction from the outset—and importantly, that they receive funding directly. As Bratteteig and Wagner point out, the “skill composition” of a design team significantly shapes how problems and solutions are defined (Bratteteig and Wagner, 2016). Including community members in core project roles enables them to define the problem space and success criteria, and ensures their labor and expertise are properly acknowledged and compensated (Pine et al., 2020). This approach also helps redistribute power within AI4SG collaborations, supporting more equitable relationships between communities and research institutions (Erete et al., 2023). Crucially, when engaging communities throughout the project, it is important to recognize and compensate for the various forms of labor they contribute—from sharing lived experience and domain expertise to facilitating connections within their communities. Future HCI research could work towards quantifying both visible and invisible labor (see (Toxtli et al., 2021) and (Li et al., 2022)), while also acknowledging that some forms of labor in community engagement, such as emotional labor, may be difficult or impossible to measure.
@@ 5.2.4. Funders’ Support: valuing community engagement as a necessary part of AI4SG work
Funders can better support community engagement in AI4SG by recognizing that relationship-building is a foundational phase of the work, which often occurs before a clear project direction is even established (Carroll and Rosson, 2013; Cooper et al., 2022). This phase is essential to community-collaborative approaches (Cooper et al., 2022), and it requires the routine presence of project teams, which takes time and emotional labor (Dearden and Tucker, 2016). Long-term relationships enable shared agendas to emerge, with community priorities guiding technology development from the outset (Dearden and Tucker, 2016; Carroll and Rosson, 2013). One concrete step funders can take is to extend the typical AI4SG funding timeline (in our corpus: with a median of one and a half years) to accommodate this early-stage work. Relationship-building alone can take months (Erete et al., 2023) or even years (Carroll and Rosson, 2013; Dearden and Tucker, 2016), and should be supported as a formal project phase, even prior to defining problems or solutions. One key challenge is determining appropriate timelines and support for this work. Future research in CCA could analyze researchers’ reflections (e.g., (Erete et al., 2023; Dearden and Tucker, 2016; Carroll and Rosson, 2013)) to better understand the duration and nature of relationship-building across projects.
@@
Furthermore, funders should also treat community engagement as a critical skill that requires training and support, on par with technical expertise. Many AI practitioners are trained in Computer Science departments that don’t routinely provide training in community engagement. As Okolo and Lin observe, AI4SG practitioners often use a hodgepodge of ad-hoc approaches to engage with end users (Okolo and Lin, 2024). Funders should recognize this knowledge gap and provide clear guidance to project teams on how to properly identify stakeholders and engage with end beneficiaries. There are multiple places where this guidance could be provided: as part of the funding call itself or additional training support that many funders already provide. Funders could consider citing existing guides in community engagement from the HCI community, such as the Equity-Centered Community Design Field Guide (Creative Reaction Lab, 2024), the Atlanta Community Engagement Playbook (Taranji Alvarado et al., 2024), and “readiness for partnership” checks (Pine et al., 2020). Other design toolkits like the Envisioning Cards (Friedman et al., 2024) might also be useful for AI4SG project teams to identify and engage with stakeholders. Since AI introduces additional hurdles to community-based designs, including requiring intensive data work (Miceli and Posada, 2022; Passi and Sengers, 2020; Sambasivan et al., 2021) and transparency issues (Okolo and Lin, 2024), HCI researchers could support funders by tailoring existing toolkits to the specific contexts of AI4SG, building on emerging Participatory AI4SG frameworks like the PACT framework (Bondi et al., 2021).
@@ 5.3. Calling for Meaningful Engagement: Recommendations for the Funding Design Process
Beyond the design of individual funding calls, we offer recommendations for funders and the HCI community on the design process of funding programs. The design process of grantmaking is influenced by funders’ connections to other stakeholders (Saha et al., 2022a). From our dataset, we observed that private tech companies were not only providers of AI products, resources, and training support, but also funders themselves. Through investment in AI4SG, private tech companies could not only expand their market influence (Henriksen and Richey, 2022; Lukacz, 2024; Nost and Colven, 2022) but also influence other funders’ beliefs about AI’s potential and their funding priorities. At the same time, the influence of communities on funding documents remains scant (c.f.  (Saha et al., 2022a)). To reorient AI4SG towards a more balanced approach, we see the need for funders to engage with intended communities and the HCI community (Section 5.3.1). We further outline implications for HCI researchers and practitioners to support the AI4SG funding design process in Section 5.3.2.
5.3.1. For AI4SG funders: engage with intended communities
Engaging with communities in the design of funding calls would enable funders to align their agendas with the needs of those communities early in the decision-making process, before projects even begin (Saha et al., 2022a; Balestrini et al., 2014). For funding programs that target projects in a region far from the funding organizations’ locations, it is essential to have regional partners with a more solid contextual understanding, such as regional research hubs. They can serve as “intermediaries” between funding organizations and intended end beneficiaries that funding organizations don’t have direct access to (Dell and Kumar, 2016; Saha et al., 2022a).
@@ 5.3.1. For AI4SG funders: engage with intended communities
5.3.2. For the HCI community: implications for practice and research
Our findings signal that funders are already engaging with Responsible AI guidelines and research, as evident in some funding calls that referenced academic works and guidebooks (Section 4.4.2). This indicates opportunities for HCI research and practical tools to inform funders’ design process. Following up on Saha et al.’s call on HCI researchers and practitioners to serve as “mediators” between decision-makers and communities’ voices (Saha et al., 2022b), we see opportunities in the AI4SG space where the HCI community can mediate between communities, funders, and other stakeholders like funders’ regional partners.
For HCI practitioners
While there is a rich body of resources for community-collaborative approaches (CCA) (Cooper et al., 2022), there exist gaps for the HCI community to bridge between frameworks and practices, such as implementing formal evaluations of participatory design projects (Bossen et al., 2016; Nguyen, 2022). One possible reason for these gaps is the lack of HCI expertise present in non-academic institutions. HCI practitioners, who work directly with non-academic stakeholders to implement frameworks, can establish partnerships with or apply for positions in funding organizations to put CCA frameworks into practice (such as consulting contracts or sabbatical positions at funding organizations like the U.S. NSF).
@@ 5.3.2. For the HCI community: implications for practice and research
Specifically, HCI practitioners can support funders in identifying key stakeholders to engage with in communities, designing training materials on community engagement for project teams, and providing community engagement workshops for both the project teams and the funders. Moreover, the HCI practitioners can help foster reflexivity in the grantmaking process, such as reflection workshops that bring funders, regional partners, representatives of intended beneficiaries, and targeted applicants together to learn from past funding program experiences, leveraging deliberation tools when considering strategies around AI (such as the Situated AI Guidebook  (Kawakami et al., 2024b)). Wellcome Trust’s design process of Mental Health Data Prizes (C8 and C9) exemplifies the benefits of feedback workshops (Section 4.4.2).
For HCI researchers
Our analyses also surface opportunities for future HCI research to further understand the role of funders in AI4SG initiatives, as well as pathways to shifting towards more balanced practices. For example, one promising direction is to interview staff members of AI4SG funding organizations to better understand the motivations, internal deliberations, and tensions between what is articulated in funding documents and the project selection process; our document analysis can serve as a basis for developing interview protocols for funders, potential applicants, and funded teams (Bowen, 2009). Future work could also engage with funded project teams to better understand their funding application strategies and needs (c.f., (Foster et al., 2024; Baxter et al., 2016)). Linking funding documents and funded project outcomes could be fruitful in identifying alignments or gaps between visions communicated through funding rhetoric and practices on the ground. Future work could also categorize commonalities of success and failure specific to AI4SG initiatives (similar to Dodson et al. (2012)’s analysis of ICTD projects). Since publicly available written materials may be biased toward success stories, focusing on one or a few case studies could yield insightful analysis. Last but not least, future work could expand the analysis to AI4SG funding documents published in languages other than English to identify funders located outside of Western societies and examine their funding agendas.
@@
Furthermore, the HCI research community can offer counter-narratives to the more techno-centric imaginaries of AI in funding calls, such as those that emphasize diverse ways of knowing and sustainable development of AI within planetary boundaries. Some emerging critical reassessments of current AI design practices and public discourse around AI include work on AI abandonment (Johnson et al., 2024), post-capitalist AI systems based on feminist values (Varon et al., 2024), Indigenous thoughts (Ofosu-Asare, 2024; Lewis et al., 2024), post-colonial computing (Sultana and Ahmed, 2019), and post-growth philosophies (Sharma et al., 2023; Kuijer and Laschke, 2024). HCI scholars can make these perspectives more accessible to funders and AI4SG stakeholders beyond the research community by translating them into public-facing texts, such as articles in ACM Interactions magazine and online posts. These textual instruments could help bridge academic insights and real-world decision-making.
@@ 6. Conclusion
Artificial Intelligence for Social Good (AI4SG) emerges as a growing body of research and practice that explores the potential of AI and data-driven technologies to tackle social problems like public health (Shi et al., 2020; Perrault et al., 2020; Tomašev et al., 2020; Rolnick et al., 2022). A growing body of work shows that AI4SG initiatives face various challenges in real-world contexts, including a lack of community engagement and project sustainability (Lin et al., 2024; Aula and Bowles, 2023; Ben Green, 2019). While prior work has focused on the roles that project teams—typically AI developers and their nonprofit partners–play in project outcomes, an emerging body of work points to the significant influence of funding agendas in shaping the direction, design, and priorities of technology initiatives aimed at social impact (Lin et al., 2024; Saha et al., 2022a; Gardner et al., 2022).
In this study, we begin addressing this gap and “study up” AI4SG funding documents, examining them through the lens of existing literature on “tech for good” projects. Our analysis reveals dissonances between AI4SG funders’ intentions for positive, sustainable social impact and the techno-centric approaches that some funding programs promoted. At the same time, we identified several more balanced approaches as well as existing gaps in funding calls. Building on our analysis, we offer recommendations for funders on funding call designs, while emphasizing the importance of community engagement in the funding call design process. Finally, we specify the roles that HCI practitioners and researchers can play to mediate between AI4SG stakeholders, supporting meaningful engagement with communities in the early stages of AI4SG projects—when funding programs are being designed.
@@ Appendix A Funding documents
C6 Data for Social Sustainability call for proposals 2021 lele{} \$80,000 (avg) + AWS credits ($\leq{}$ \$20,000 avg) 1 year Amazon Private Tech Company United States Global human rights protections “full-time faculty member or permanent NGO researcher” (Amazon, 2021)